OpenStudy (anonymous):

WILL GIVE MEDAL AND FAN!! Which fractions are greater than -4/7? Choose exactly TWO answers that are correct. A -1/2 B -5/8 C -2/3 D -6/11 I HAVE NO CLUE HOW TO DO THIS. PLEASE HELP.

OpenStudy (welshfella):

one way to do this would be to convert each fraction to a decimal so -4/7 = -0.5714 -1/2 = -0.5 -5/8 = -2.625 can you continue? Correct to 4 decimal places if necessary

OpenStudy (welshfella):

sorry the last one -5/8 = -0.625

OpenStudy (anonymous):

-2/3 would be what?

OpenStudy (welshfella):

use your calculator :- Divide -2 by 3

OpenStudy (anonymous):

0.6??

OpenStudy (welshfella):

no it would be a negative just use your calculator -2 / 3 and press exe

OpenStudy (welshfella):

or divide might be 

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

theres no symbol like that theres * / - +

OpenStudy (welshfella):

use /

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

its not working.

OpenStudy (anonymous):

im getting the same results over and over again

OpenStudy (welshfella):

what result is that?

OpenStudy (anonymous):

0.6

OpenStudy (anonymous):

or -0.6

OpenStudy (anonymous):

each time.

OpenStudy (welshfella):

well thats not a very good calculator have you tried the one on your PC Are you on a PC?

OpenStudy (anonymous):

i tried both.

OpenStudy (anonymous):

on PC i get -0.666 etc.

OpenStudy (welshfella):

thats it round it to 4 decimals = -0.6667

OpenStudy (welshfella):

now enter -6 / 11 =

OpenStudy (anonymous):

-0.545

OpenStudy (welshfella):

right -0.5455 to 4 decimals now its easy just write them all down and you can find the 2 which are greater that -4/7

OpenStudy (welshfella):

so -4/7 = -0.5714 -1/2 = -0.5 -5/8 = -2.625 -2/3 = -0.6667 -6/1 = -0.5455

OpenStudy (welshfella):

thats -0.625 for -5/8

OpenStudy (anonymous):

-2/3 and -5/8?

OpenStudy (welshfella):

No note they are negative numbers -0.4 is greater than -0.5

OpenStudy (welshfella):